OpenStudy (anonymous):

Find the slope-intercept equation of the line that passes through the point (-2, 4) and is parallel to y = -3x + 6.

OpenStudy (anonymous):

Please help me!

OpenStudy (anonymous):

ok... when they tell u PARALLEL means that both line have the same slope

OpenStudy (anonymous):

the slopes of parallel lines are equal.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so, in the eq. that they r giving u y=-3x+6

OpenStudy (anonymous):

can u identify the slope?

OpenStudy (anonymous):

-3?

OpenStudy (anonymous):

perfect

OpenStudy (anonymous):

so there is a formula to find the slope intercep.. y-y1=m(x-x1)

OpenStudy (anonymous):

y-4=-3(x-(-2))

OpenStudy (anonymous):

and that is like the one we just did

OpenStudy (anonymous):

okay, so it's y=3x+10, then what?

OpenStudy (anonymous):

y-4=-3(x+2) y-4=-3x-6 y=-3x-2

OpenStudy (anonymous):

chek ur signs

OpenStudy (anonymous):

oh, opps... I forgot the -...

OpenStudy (anonymous):

cool

OpenStudy (anonymous):

okay, thanks!

OpenStudy (anonymous):

from second equation we know that slope is -3 Point/Slope Form of an Equation A line going through the point (x1,y1) and having slope of m would have the equation y-y1 = m (x - x1) This gives y-4=-3[x-(-2)] y-4 = -3[x + 2] y - 4 = -3x -6 y = -3x -6 + 4 y = -3x - 2

OpenStudy (anonymous):

thanks what we got,, thanks for the summary